OpenStudy (anonymous):

twice the difference of a number and 3 is equal to three times the sum of the number and 2. find the number

OpenStudy (anonymous):

Solve the following for x: 2 (x - 3) = 3 (x + 2)

OpenStudy (anonymous):

i come up with 5 ???

OpenStudy (anonymous):

-6 + 2 x = 6 + 3 x -12 = x